# On the effectiveness of discrete representations in sparse mixture of experts

### v1

@@ Abstract
Sparse mixture of experts (SMoE) is an effective solution for scaling up model capacity without increasing the computational costs. A crucial component of SMoE is the router, responsible for directing the input to relevant experts; however, it also presents a major weakness, leading to routing inconsistencies and representation collapse issues. Instead of fixing the router like previous works, we propose an alternative that assigns experts to input via indirection, which employs the discrete representation of input that points to the expert. The discrete representations are learnt via vector quantization, resulting in a new architecture dubbed Vector-Quantized Mixture of Experts (VQMoE). We provide theoretical support and empirical evidence demonstrating the VQMoE’s ability to overcome the challenges present in traditional routers. Through extensive evaluations on both large language models and vision tasks for pre-training and fine-tuning, we show that VQMoE achieves a 28% improvement in robustness compared to other SMoE routing methods, while maintaining strong performance in fine-tuning tasks.
@@ 1 Introduction
Scaling Transformers with data and compute has demonstrated unprecedented successes across various domains such as natural language processing (NLP) tasks (Du et al., 2022; Fedus et al., 2022; Zhou et al., 2024), and visual representation learning (Riquelme et al., 2021a; Shen et al., 2023b). However, training and inference of a single large Transformer-based model might require hundreds of thousands of compute hours, costing millions of dollars (Kaddour et al., 2023). This issue has motivated contemporary studies to investigate Sparse Mixture-of-Experts (SMoE) (Shazeer et al., 2017; Zoph et al., 2022; Xue et al., 2024; Jiang et al., 2024). SMoE models that are inspired by (Jacobs et al., 1991a) usually include a set of experts sharing the same architecture and a router that activates only one or a few experts for each input. Compared to dense models of the same size, SMoE counterparts significantly reduce inference time thanks to not using all experts simultaneously (Artetxe et al., 2022; Krajewski et al., 2024).
However, training SMoEs remains a challenge due to representation collapse, that is, either a small number of experts receive most of the routed tokens or all experts converge to learn similar representations. To tackle the issue, several works (Chi et al., 2022; Chen et al., 2023a; Do et al., 2023) have focused on router policy improvement. However, these do not touch a fundamental question, ‘Do we really need a router in the first place?’ Our research suggests that adopting a discrete representation could help solve the challenges currently faced by the router method. Discrete representation learning in the context of SMoE is motivated by its ability to capture structured and interpretable patterns within data, aligning with the way that humans categorize and process information through distinct symbols, like tokens. This approach enables better generalization and facilitates knowledge transfer across different contexts. Additionally, discrete representations provide a robust and efficient mechanism for selecting and routing inputs to the appropriate experts by clustering them more effectively. By bridging the gap between discrete and continuous representations, this method leads to more stable and interpretable expert assignments, helping to mitigate issues such as representation collapse and overfitting, which are common challenges in SMoE training.
@@
In summary, the primary contributions of this paper are threefold: (1) we theoretically demonstrate that learning a discrete representation is an optimal approach for expert selection and that VQMoE inherently addresses the issue of representation collapse; (2) we propose the use of the Vector Quantization method to learn cluster structures and resolve related challenges; and (3) we conduct extensive experiments on large language models and vision pre-training and fine-tuning tasks, providing an in-depth analysis of VQMoE’s behavior to showcase its effectiveness.
@@ 2 Related Work
Discrete Representation.  Discrete representations align well with human thought processes; for example, language can be understood as a series of distinct symbols. Nevertheless, the use of discrete variables in deep learning has proven challenging, as evidenced by the widespread preference for continuous latent variables in most current research. VQVAE (van den Oord et al., 2017) implements discrete representation in Variational AutoEncoder (VAE) (Kingma & Welling, 2022) using vector quantisation (VQ). IMSAT (Hu et al., 2017) attains a discrete representation by maximizing the information-theoretic dependency between data and their predicted discrete representations. Recent works follow up the vector quantisation ideas and make some enhancements for VAE, for example: (Yu et al., 2022);  (Mentzer et al., 2023); and (Yang et al., 2023). Mao et al. (2022) utilize a discrete representation to strengthen Vision Transformer (ViT) (Dosovitskiy et al., 2021). To the best of our knowledge, our paper is the first to learn a discrete representation of Sparse Mixture of Experts.
@@ 3.1 Preliminaries
Sparse Mixture of Experts.  Sparse Mixture of Experts (SMoE) is often a transformer architecture that replaces the MLP layers in standard transformers with Mixture of Experts (MoE) layers (Shazeer et al., 2017). Given $\mathbf{x}\in\mathbb{R}^{n\times d}$ as the output of the multi-head attentions (MHA), the output of SMoE with $N$ experts is a weighted sum of each expert’s computation $E_{i}(x)$ by the router function $\mathcal{S}(x)$:
$f_{\mathrm{SMoE}}(\bm{x})=\sum_{i=1}^{N}\mathcal{S}(\bm{x})_{i}\cdot E_{i}(\bm{x})=\sum_{i=1}^{N}\mathcal{S}(\bm{x})_{i}\cdot\bm{W}_{\mathrm{FFN}_{i}}^{2}\phi\left(\bm{W}_{\mathrm{FFN}_{i}}^{1}\bm{x}\right)$ (1)
Where $\mathcal{S}(x)$ is computed by $TopK$ function as equation (2) that determines the contribution of each expert to the SMoE output.
$\mathcal{S}(\bm{x})=\operatorname{TopK}(\operatorname{softmax}(\mathcal{G}(\bm{x})),k);\operatorname{TopK}(\bm{v},k)=\begin{cases}\bm{v_{i}}&\text{ if }\bm{v_{i}}\text{ is in the top $k$ largest of }\bm{v}\\ -\infty&\text{ otherwise }\end{cases}$ (2)
Discrete Representation Learning.van den Oord et al. (2017) propose VQVAE, which uses Vector Quantisation (VQ) to learn a discrete representation. Given an input $x\in\mathbb{R}^{n\times d}$, VQVAE discretized the input into a codebook $V\in\mathbb{R}^{K\times d}$ where $K$ is the codebook size and $d$ is the dimension of the embedding. Let denote $z_{v}(x)\in\mathbb{R}^{n\times d}$ denotes the output of the VQVAE and $\mathbf{1}()$ is the indicator function. The discrete representation $z_{q}(x_{i})=v_{k},\quad\text{ where }\quad k=\operatorname{argmin}_{j}\left\|z_{v}(x_{i})-v_{j}\right\|_{2}$ is achieved by vector quantizer $q_{\theta}$ that maps an integer $z$ for each input $x$ as:
@@ 3.2 Vector-Quantized Mixture of Experts (VQMoE)
Pre-training VQMoE.  Existing Sparse Mixture of Experts (SMoE) models learn continuous representations and select experts based on routing scores derived from token-expert embeddings. In this paper, we propose a novel architecture that learns simultaneously continuous and discrete representations at a training phase as Figure 1(a). The continuous representation enables the model to capture complex structures in the data, while the discrete representation learns latent representation from data and then transfers the knowledge to downstream tasks. Given $\mathbf{x}\in\mathbb{R}^{n\times d}$ as the output of the MHA and $f^{\mathrm{v}}$ is a vector quantization operator, the output of the VQMOE layer at the Pre-training phase as follows:
$f^{\mathrm{VQMoE}}(\bm{x})=g\left(x\right)_{c}f^{\mathrm{SMoE}}(\bm{x})+g\left(x\right)_{d}\sum_{l=1}^{K}f_{l}^{\mathrm{FFN}}(\bm{\tilde{x_{l}}}),$ (4)
Where $\tilde{x_{l}}=v_{k}$ if $x_{l}\in V_{l}$ codebook, otherwise $\tilde{x_{l}}=\vec{0}$ ; $f_{l}^{\mathrm{FFN}}({\tilde{x_{l}}})$ corresponds to the expert associated with the $V_{l}$ codebook; $g(x)_{c}(x)=col_{0}(G(x))$, $g(x)_{d}(x)=col_{1}(G(x))$ is gating function for continuous and discrete representation with $G(x)=\operatorname{softmax}(W_{g}^{T}\times x)$. $W_{g}^{T}\in\mathbb{R}^{2\times d}$ is a learnable weight and $K$ is number of codes.
Fine-tuning VQMoE. According to  (Geva et al., 2021), the Feed-forward layers (FFN) constitute two-thirds of a transformer model’s parameters. Thus, VQMoE enhances the robustness and efficiency of the Mixture of Experts by leveraging the discrete representations learned during the Pre-training phase. For further details, the output of VQMoE during the fine-tuning stages only requires the discrete representation part as Figure  1(b), leading to the following output from the VQMoE layer in the fine-tuning phase:
$f^{\mathrm{VQMoE}}(\bm{x})=\sum_{l=1}^{K}f_{l}^{\mathrm{FFN}}(\bm{\tilde{x}_{l}})$ (5)
@@ 3.3 Training Procedure
$L=\mathcal{L}_{\text{task }}+\alpha(\left\|\operatorname{sg}\left[z_{v}(x)\right]-v\right\|_{2}^{2}+\beta\left\|z_{v}(x)-\operatorname{sg}[v]\right\|_{2}^{2})$ (6)
@@
Fine-tuning.  For downstream tasks, we fine-tune the pretraining model by utilizing the codebook learned from the Equation  6 by freezing all parameters at the Vector Quantization module. Thus, the training objective simply becomes: $L=\mathcal{L}_{\text{task }}$.
@@
4 Theoretical Guarantees of VQMoE
4.1 Theory Analysis
Problem settings.  We consider an MoE layer with each expert being an MLP layer which is trained by gradient descent and input data $\left\{\left(\mathbf{x}_{i},y_{i}\right)\right\}_{i=1}^{n}$ generated from a data distribution $\mathcal{D}$. Same as  (Chen et al., 2022);  (Dikkala et al., 2023), we assume that the MoE input exhibits cluster properties, meaning the data is generated from $K$ distinct clusters $(C_{1},C_{2},...,C_{k})$.
Inspired by  (Dikkala et al., 2023), we conceptualize the router in Sparse Mixture of Experts as a clustering problem. This leads us to define a consistent router in Definition  B.1. Furthermore, we introduce a definition for an inconsistent router in SMoE as outlined in Definition  B.2, along with the concept of inconsistent expert selection presented in Theorem 4.1 during the training of SMoE.
@@ 4.1 Theory Analysis
Let $f_{MHA}$ be a multi-head attention (MHA) function producing an output $x\in\mathbb{R}^{n\times d}$, and consider $N$ experts with embeddings $e_{i}$ for expert $i$ where $i\in[1,N]$. Assume that $f_{MHA}$ converges at step $t_{m}$, while the expert embeddings $e$ converge at step $t_{e}$, with $t_{m}\gg t_{e}$. For each output $x$, the expert $K\in[1,N]$ is selected such that
$K=\arg\min_{j\in[1,N]}\text{dist}(x,e_{j}).$
@@
The proof of Theorem  4.1 is given in Appendix  A, and we have the following insights. Theorem  4.1 implies that an expert selection process by a router as the conventional SMoE leads to the inconsistent router. Indeed, the router layer is designed as a simple linear layer, $x$ is the output of MHA function in practice. In practice, an SMoE router is significantly simpler than the MHA function. Consequently, this design leads to the router functioning as an inconsistent router, contributing to the representation collapse issue and instability during training.
@@
Given input data partitioned into $k$ clusters $(C_{1},C_{2},\ldots,C_{k})$ and a mixture of experts (MoE) layer with $k$ experts $(E_{1},E_{2},\ldots,E_{k})$, the assignment of each cluster $C_{i}$ to expert $E_{i}$ for $i\in[1,k]$ constitutes an optimal expert selection solution.
@@
Proposition  4.2 demonstrates that if we are given a clustering structure as input, assigning each part of the input to its corresponding expert results in an optimal expert selection. This implies that learning a discrete representation and directing each component to the appropriate expert yields an optimal solution. The proof of Proposition  4.2 can be found in Appendix  A.
4.2 VQMoE solves Representation Collapse by Design
The representation collapse problems in SMoE, which leads all experts to learn the same thing, first declared by  (Chi et al., 2022). Same as  (Chi et al., 2022);  (Do et al., 2023), we illustrate the presentation collapse issue by Jacobian matrix approach. Indeed, Jacobian matrix of SMoE with respect to $x\in\mathbb{R}^{n\times d}$ is followed as:
$\bm{J_{SMoE}}=\mathcal{S}(x)_{k}\bm{J}^{\mathrm{FFN}}+\sum_{j=1}^{N}\mathcal{S}(x)_{k}\left(\delta_{kj}-S_{j}\right)\bm{E}(x)_{i}\bm{e}_{j}^{\top}=\mathcal{S}(x)_{k}\bm{J}^{\mathrm{FFN}}+\sum_{j=1}^{N}\bm{c}_{j}\bm{e}_{j}^{\top},$ (8)
where $\bm{c}_{j}=\mathcal{S}(x)_{k}\left(\delta_{kj}-S_{j}\right)\bm{E}(x)_{i}$. Equation  8 consists two terms: (1) $\mathcal{S}(x)_{k}\bm{J}^{\mathrm{FFN}}$ represents a contribution from input token and experts to the final output; (2) $\sum_{j=1}^{N}\bm{c}_{j}\bm{e}_{j}^{\top}$ indicates to learn better gating function to minimize the task loss. Moreover, Equation  8 is suggested to be updated toward a linear combination of the expert embeddings. Since $N<<d$ in practice, the above equation shows representation collapse from $\mathbb{R}^{d}$ to $\mathbb{R}^{N}$.
Compared to SMoE, does VQMoE reduce the representation collapse issue? To answer the essential question, we calculate the Jacobian matrix of VQMoE with respect to $x\in\mathbb{R}^{n\times d}$ is given by:
$\bm{J_{VQMoE}}=g\left(x\right)_{c}J_{SMoE}+J_{g\left(x\right)_{c}}f_{\mathrm{SMoE}}(x)+g\left(x\right)_{d}J_{VQ}+J_{g\left(x\right)_{d}}f_{\mathrm{VQMoE}}(x)$ (9)
Equation  9 is written shortly as below:
$\bm{J_{VQMoE}}=J_{1}+\sum_{j=1}^{N}c_{j}e_{j}^{\top}+\sum_{l=1}^{K}d_{l}e_{l}^{\top}+\sum_{m\in{c,d}}g_{m}e_{m}^{\top}=J_{1}+\sum_{j=1}^{N+K+2}o_{j}e_{j}^{\top}$ (10)
where $J_{1}=\mathcal{S}(x)_{k}\bm{J}^{\mathrm{FFN}}$ ; $c_{j}=\mathcal{S}(x)_{k}\left(\delta_{kj}-S_{j}\right)\bm{E}(x)_{i}$ ; $d_{l}=g\left(x\right)_{d}$ (due to the vector quantization operator using pass gradient trick (van den Oord et al., 2017)); $g_{m}=\mathcal{S}(x)_{m}\left(\delta_{mk}-S_{k}\right)f_{m}\textit{ where }f_{m}\in[f_{\mathrm{SMoE}}(x),f_{\mathrm{VQMoE}}]$.

### v2

# On the Role of Discrete Representation in Sparse Mixture of Experts

@@ Abstract
Sparse Mixture of Experts (SMoE) is an effective solution for scaling up model capacity without increasing the computational costs. A crucial component of SMoE is the router, responsible for directing the input to relevant experts; however, it also presents a major weakness, leading to routing inconsistencies and representation collapse issues. Instead of fixing the router like previous works, we propose an alternative that assigns experts to input via indirection, which employs the discrete representation of input that points to the expert. The discrete representations are learned via vector quantization, resulting in a new architecture dubbed Vector-Quantized Mixture of Experts (VQMoE). We provide theoretical support and empirical evidence demonstrating the VQMoE’s ability to overcome the challenges present in traditional routers. Through extensive evaluations on both large language models and vision tasks for pre-training and fine-tuning, we show that VQMoE achieves a 28% improvement in robustness compared to other SMoE routing methods while maintaining strong performance in fine-tuning tasks.
@@ 1 Introduction
Scaling Transformer models with increasing data and computational resources has led to remarkable advances across a wide range of domains, including natural language processing (NLP) (Du et al., 2022; Fedus et al., 2022; Zhou et al., 2024) and visual representation learning (Riquelme et al., 2021a; Shen et al., 2023b). Despite these successes, training and deploying large-scale dense Transformer models often require substantial computational resources, frequently amounting to hundreds of thousands of GPU hours and incurring costs in the millions of dollars (Kaddour et al., 2023). To address this scalability bottleneck, Sparse Mixture of Experts (SMoE) architectures have emerged as a promising alternative (Shazeer et al., 2017; Zoph et al., 2022; Xue et al., 2024; Jiang et al., 2024). Inspired by classical Mixture of Experts formulations (Jacobs et al., 1991a), SMoE models consist of multiple expert subnetworks with shared architectures, where a routing mechanism dynamically selects a small subset of experts (often one or two) for each input token. This sparsity significantly reduces inference costs compared to dense counterparts of similar model capacity (Artetxe et al., 2022; Krajewski et al., 2024), making SMoEs attractive for efficient scaling.
Despite their efficiency benefits, SMoEs face critical training challenges, most notably, representation collapse. This phenomenon occurs when only a small subset of experts are frequently activated, or when all experts converge to similar representations, thereby negating the diversity and specialization that the architecture is intended to promote. Prior works have sought to mitigate this issue by improving the routing policy through regularization and auxiliary losses (Chi et al., 2022; Chen et al., 2023a; Do et al., 2023). However, these approaches focus on the routers improvement rather than questioning its necessity.
In this work, we explore a more fundamental question: Is an explicit router necessary at all? We argue that incorporating discrete representations offers a principled alternative. Discrete latent variables are inherently suited to capturing structured and interpretable patterns within data, aligning with the symbolic nature of human cognition, where concepts are often discretized as words, tokens, or categories. In the SMoE context, discrete representations can improve input routing by naturally clustering similar inputs, thereby enhancing expert specialization and utilization without relying solely on a learned gating mechanism.
@@
In summary, the primary contributions of this paper are as follows:
  • We theoretically demonstrate that learning discrete representations provides an effective mechanism for expert selection, and that VQMoE intrinsically mitigates the problem of representation collapse.
  • We propose the use of vector quantization (VQ) to learn structured and interpretable expert clusters.
  • We conduct extensive experiments on large language models as well as vision pre-training and fine-tuning tasks to validate the effectiveness of our method.
  • We provide a comprehensive analysis of VQMoE’s behavior, offering insights into its performance and robustness.
@@ 2 Related Work
Discrete Representation.  Discrete representations align well with human thought processes; for example, language can be understood as a series of distinct symbols. Nevertheless, the use of discrete variables in deep learning has proven challenging, as evidenced by the widespread preference for continuous latent variables in most current research. VQVAE (van den Oord et al., 2017) implements discrete representation in Variational AutoEncoder (VAE) (Kingma & Welling, 2022) using vector quantization (VQ). IMSAT (Hu et al., 2017) attains a discrete representation by maximizing the information-theoretic dependency between data and their predicted discrete representations. Recent works follow up the vector quantization ideas and make some enhancements for VAE, for example: (Yu et al., 2022);  (Mentzer et al., 2023); and (Yang et al., 2023). Mao et al. (2022) utilize a discrete representation to strengthen Vision Transformer (ViT) (Dosovitskiy et al., 2021). To the best of our knowledge, our paper is the first to learn a discrete representation of Sparse Mixture of Experts.
@@ 3.1 Preliminaries
Sparse Mixture of Experts.  Sparse Mixture of Experts (SMoE) is a variant of the transformer architecture in which the conventional feed-forward layers (MLPs) are replaced with Mixture of Experts (MoE) layers (Shazeer et al., 2017). Given an input $\bm{x}\in\mathbb{R}^{n\times d}$, which represents the output of the multi-head attention (MHA) module, the SMoE layer computes a sparse weighted combination over a set of $N$ expert networks. Each expert is typically a feed-forward neural network $FFN_{i}(\bm{x})$, and its contribution to the final output is determined by a routing function $\mathcal{S}(\bm{x})$. The resulting output of the SMoE layer is given by:
$f^{\mathrm{SMoE}}(\bm{x})$ $=\sum_{i=1}^{N}\mathcal{S}(\bm{x})_{i}\cdot FFN_{i}(\bm{x})$ (1)
$=\sum_{i=1}^{N}\mathcal{S}(\bm{x})_{i}\cdot\bm{W}_{\mathrm{FFN}_{i}}^{2}\,\phi\left(\bm{W}_{\mathrm{FFN}_{i}}^{1}\bm{x}\right),$
[…]
$\mathcal{S}(\bm{x})$ $=\operatorname{TopK}(\operatorname{softmax}(\bm{W}_{e}\bm{x}),k),$ (2)
[…]
This sparse selection mechanism ensures that only a small subset of experts are activated for each input, which significantly reduces computational cost while retaining model capacity.
Discrete Representation Learning.van den Oord et al. (2017) propose VQVAE, which uses Vector Quantization (VQ) to learn a discrete representation. Given an input $x\in\mathbb{R}^{n\times d}$, VQVAE discretized the input into a codebook $V\in\mathbb{R}^{K\times d}$ where $K$ is the codebook size and $d$ is the dimension of the embedding. Let denote $z_{v}(x)\in\mathbb{R}^{n\times d}$ the output of the VQVAE and $\mathbf{1}()$ is the indicator function. The discrete representation $z_{q}(x_{i})=v_{k},\quad\text{ where }\quad k=\operatorname{argmin}_{j}\left\|z_{v}(x_{i})-v_{j}\right\|_{2}$ is achieved by vector quantizer $q_{\theta}$ that maps an integer $z$ for each input $x$ as:
@@ 3.2 Vector-Quantized Mixture of Experts (VQMoE)
Pre-training VQMoE.  Traditional Sparse Mixture of Experts (SMoE) models utilize continuous token representations and route them to experts based on learned token-expert affinity scores. We propose a novel architecture, VQMoE, that learns both continuous and discrete representations jointly during pre-training (see Figure 1(a)). The continuous component captures fine-grained data patterns, while the discrete component, learned via vector quantization, encodes robust latent structure useful for downstream transfer.
Let $\bm{x}\in\mathbb{R}^{n\times d}$ denote the input to the VQMoE layer (e.g., output from a multi-head attention block), and let $f^{\mathrm{vq}}$ denote the vector quantization operator. The VQMoE output during pre-training is defined as:
$f^{\mathrm{VQMoE}}(\bm{x})=\underbrace{g_{c}(\bm{x})\cdot f^{\mathrm{SMoE}}(\bm{x})}+\underbrace{g_{d}(\bm{x})\cdot\sum_{l=1}^{K}f_{l}^{\mathrm{FFN}}(\tilde{\bm{x}}_{l})}$ (4)
(Continuous representation)     (Discrete representation)
In this formulation, $f^{\mathrm{SMoE}}(\bm{x})$ denotes the output from a standard Sparse Mixture of Experts (SMoE) layer, capturing the continuous expert representations. The second term corresponds to the discrete representation, where each $f_{l}^{\mathrm{FFN}}$ is the $l$-th feedforward expert network. The input to each discrete expert, denoted as $\tilde{\bm{x}}_{l}$, is determined by vector quantization: specifically, $\tilde{\bm{x}}_{l}=\bm{v}_{k}$ if the input vector $\bm{x}_{l}$ is assigned to the $l$-th codebook vector $\bm{v}_{k}$; otherwise, it is set to the zero vector, i.e., $\tilde{\bm{x}}_{l}=\bm{0}$. Here, $K$ is the number of vector quantization codebooks, and $\bm{v}_{k}$ is a learned codebook vector assigned by $f^{\mathrm{vq}}$. The gating functions $g_{c}(\bm{x})$ and $g_{d}(\bm{x})$ as Equation  5, modulate the contributions of the continuous and discrete pathways, respectively, and are typically computed based on the input $\bm{x}$ through learnable mechanisms.
$\begin{bmatrix}g_{c}(\bm{x})\ g_{d}(\bm{x})\end{bmatrix}=\operatorname{softmax}(W_{g}\bm{x}),\quad W_{g}\in\mathbb{R}^{2\times d}$ (5)
To address the mismatch between the number of codebook vectors and the number of expert networks, we introduce a flexible code strategy. This approach enables consistent routing from quantized representations to experts, even when the two quantities differ. Specifically, we define a hash-based mapping using a modulo operation. Let $i_{\mathrm{cb}}$ denote the index of a codebook vector, and let $i_{\mathrm{exp}}$ denote the index of the corresponding expert. The mapping is given by:
$i_{\mathrm{exp}}=i_{\mathrm{cb}}\bmod N,$ (6)
where $N$ is the total number of experts. This ensures each codebook index is deterministically assigned to one of the available experts
Fine-tuning VQMoE.  Based on insights from Geva et al. (2021), which note that feed-forward layers (FFNs) constitute a significant portion of a transformer’s parameters, we adopt a lightweight fine-tuning strategy that retains only the discrete path of the VQMoE. This allows efficient adaptation while leveraging pre-trained latent representations (see Figure 1(b)). The fine-tuning output becomes:
$f^{\mathrm{VQMoE}}(\bm{x})=\sum_{l=1}^{K}f_{l}^{\mathrm{FFN}}(\tilde{\bm{x}}_{l})$ (7)
@@ 3.3 Training Procedure
$L=\mathcal{L}^{\text{task }}+\alpha(\left\|\operatorname{sg}\left[z_{v}(x)\right]-v\right\|_{2}^{2}+\beta\left\|z_{v}(x)-\operatorname{sg}[v]\right\|_{2}^{2})$ (8)
@@
In Equation 8, $\mathcal{L}^{\text{task}}$ denotes the task-specific loss, which depends on applications. For example, in language modeling tasks, $\mathcal{L}^{\text{task}}$ is typically defined as the negative log-likelihood (NLL) of the target tokens (Dai et al., 2019c), promoting accurate next-token prediction. In image classification tasks, $\mathcal{L}^{\text{task}}$ is usually implemented as the cross-entropy loss between the predicted class distribution and the ground-truth label (He et al., 2015), encouraging correct class assignment.
Fine-tuning.  For downstream tasks, we fine-tune the pretraining model by utilizing the codebook learned from the Equation  8 by freezing all parameters at the Vector Quantization module. Thus, the training objective simply becomes: $L=\mathcal{L}^{\text{task }}$.
@@
4 Theory Analysis
4.1 Optimal Experts Selection
Problem settings.  We consider an MoE layer with each expert being an MLP layer which is trained by gradient descent and input data $\left\{\left(\mathbf{x}_{i},y_{i}\right)\right\}_{i=1}^{n}$ generated from a data distribution $\mathcal{D}$. Same as  (Chen et al., 2022);  (Dikkala et al., 2023), we assume that the MoE input exhibits cluster properties, meaning the data is generated from $N$ distinct clusters $(C_{1},C_{2},...,C_{N})$.
Definition 4.1 (Consistent Router)
A sequence of points $x_{1},x_{2},\ldots,x_{n}$ and a corresponding sequence of clusters $C_{1},C_{2},\ldots,C_{N}$ are said to be consistent if, for every point $x_{p}\in C_{i}$, the condition
$\text{dist}(x_{p},u_{i})\leq\min_{j\neq i}\text{dist}(x_{p},u_{j})$
Definition 4.2 (Inconsistent Router)
A sequence of points $x_{1},x_{2},\ldots,x_{n}$ and a corresponding sequence of clusters $C_{1},C_{2},\ldots,C_{N}$ are said to be inconsistent if there exists a point $x_{p}\in C_{i}$ such that
$\text{dist}(x_{p},u_{i})>\min_{j\neq i}\text{dist}(x_{p},u_{j}),$
Inspired by  (Dikkala et al., 2023), we conceptualize the router in Sparse Mixture of Experts as a clustering problem. This leads us to define a consistent router in Definition  4.1. Furthermore, we introduce a definition for an inconsistent router in SMoE as outlined in Definition  4.2, along with the concept of inconsistent expert selection presented in Theorem 4.3 during the training of SMoE.
@@ 4.1 Optimal Experts Selection
Let $f_{MHA}$ be a multi-head attention (MHA) function producing an output $x\in\mathbb{R}^{n\times d}$, and consider $N$ experts with embeddings $e_{i}$ for expert $i$ where $i\in[1,N]$. Assume that $f_{MHA}$ converges at step $t_{m}$, while the expert embeddings $e$ converge at step $t_{e}$, with $t_{m}\gg t_{e}$. For each output $x$, an expert $P\in[1,N]$ is selected such that
$P=\arg\min_{j\in[1,N]}\text{dist}(x,e_{j}).$
@@
The proof of Theorem  4.3 is given in Appendix  A.1.2, and we have the following insights. Theorem  4.3 implies that an expert selection process by a router as the conventional SMoE leads to the inconsistent router. Indeed, the router layer is designed as a simple linear layer, $x$ is the output of MHA function in practice; and an SMoE router is significantly simpler than the MHA function. Consequently, this design leads to the router functioning as an inconsistent router, contributing to the representation collapse issue and instability during training.
@@
Given input data partitioned into $N$ clusters $(C_{1},C_{2},\ldots,C_{N})$ and a mixture of experts (MoE) layer with $N$ experts $(E_{1},E_{2},\ldots,E_{N})$, the assignment of each cluster $C_{i}$ to expert $E_{i}$ for $i\in[1,k]$ constitutes an optimal expert selection solution.
@@
Proposition  4.4 demonstrates that if we are given a clustering structure as input, assigning each part of the input to its corresponding expert results in an optimal expert selection. This implies that learning a discrete representation and directing each component to the appropriate expert yields an optimal solution. The proof of Proposition  4.4 can be found in Appendix  A.1.3.
4.2 Experts Representation Collapse
The representation collapse problem in Sparse Mixture of Experts (SMoE), where all experts converge to similar representations, was first highlighted by Chi et al. (2022). Following Chi et al. (2022) and Do et al. (2023), we analyze this issue using the Jacobian matrix of the model output with respect to the input $x\in\mathbb{R}^{n\times d}$. The Jacobian for SMoE is expressed as:
$\bm{J}^{\text{SMoE}}$ $=\mathcal{S}(x)_{k}\bm{J}^{\text{FFN}}+\sum_{j=1}^{N}\mathcal{S}(x)_{k}(\delta_{kj}-\mathcal{S}(x)_{j})\bm{E}(x)_{i}\bm{e}_{j}^{\top}$ (10)
$=\mathcal{S}(x)_{k}\bm{J}^{\text{FFN}}+\sum_{j=1}^{N}\bm{c}_{j}\bm{e}_{j}^{\top},$
where $\bm{c}_{j}=\mathcal{S}(x)_{k}(\delta_{kj}-\mathcal{S}(x)_{j})\bm{E}(x)_{i}$, $\bm{J}^{\text{FFN}}$ is the Jacobian of the selected expert’s feedforward network, and $\bm{e}_{j}$ are the expert embedding vectors. Equation 10 consists of two components: $\mathcal{S}(x)_{k}\bm{J}^{\text{FFN}}$ the main signal path from the input to the output through the selected expert; and $\sum_{j=1}^{N}\bm{c}_{j}\bm{e}_{j}^{\top}$ the contribution from the gating function’s gradient with respect to the expert embeddings.
Since the summation over expert embeddings lies in a subspace of dimension $N$, and typically $N\ll d$, this projection restricts the output space from $\mathbb{R}^{d}$ to $\mathbb{R}^{N}$, which effectively causes representation collapse.
Jacobian Analysis of VQMoE.  To examine whether VQMoE mitigates this collapse, we derive the Jacobian of the VQMoE output with respect to the input $x\in\mathbb{R}^{n\times d}$. The detailed expression of the VQMoE Jacobian matrix is provided in Section A.1.1. Specifically, we have:
$\bm{J}^{\text{VQMoE}}$ $=g_{c}(\bm{x})\cdot\bm{J}^{\text{SMoE}}+\frac{\partial g_{c}(\bm{x})}{\partial\bm{x}}f^{\mathrm{SMoE}}(\bm{x})$ (11)
$\quad+g_{d}(\bm{x})\cdot\sum_{l=1}^{K}\bm{J}^{\text{FFN}}_{l}+\frac{\partial g_{d}(\bm{x})}{\partial\bm{x}}\sum_{l=1}^{K}f_{l}^{\mathrm{FFN}}(\tilde{\bm{x}}_{l})$
$=J_{1}+\sum_{j=1}^{N+K+2}o_{j}\bm{e}_{j}^{\top}.$